OpenStudy (anonymous):

Two planes are flying around the world. The first plane is flying al 300 miles per hour and is 350 miles ahead of the second plane. The second plane is flying at 400 mph. How long will take for the second plane to catch the first plane?.

OpenStudy (shane_b):

The distance between the planes is 350 miles and plane 2 is gaining on the first plane at a rate of 100 mph....so what do you think?

OpenStudy (lgbasallote):

let the initial distance of first plane = d initial distance of second plane = d - 350 speed of first plane = 300 speed of second plane = 400 final distance - initial distance = speed x time final distance = speed x time + initial distance final distance of first plane = 300t + d final distance of second plane = 400t + d - 350 when they catch up the final distance will be same so we can equate them to one another 300t + d = 400t + d - 350 400t - 300t = d - d + 350 100t = 350 t = 350/100 t = 3.5 hours make sense?

OpenStudy (anonymous):

3.5 hours= 3 hours and a half right?

OpenStudy (shane_b):

Yes.

OpenStudy (lgbasallote):

yes it is

OpenStudy (anonymous):

Thank you all so very much i hope you all go to heaven! take care!